# Superdiffusion in random two dimensional system with time-reversal symmetry and long-range hopping

### v3

@@ Abstract
This universal scaling is limited to systems with a short-range hopping, while the hopping decreasing with the distance as $V(\mathbf{r})\propto r^{-d}$ or slower leads to delocalization of all states [1; 10; 11; 12; 13; 14] except for the marginal case of diverging Fourier transform of a hopping amplitude [15; 16; 17; 18; 19; 20; 21; 22; 23; 24; 25; 26]. If disorder is strong, eigenstates of the problem with the long-range hopping $V(\mathbf{r})\propto r^{-d}$ possess a multifractal structure and the time dependent displacement of the particle $r$ obeys the law $r\propto t^{1/d}$, which is subdiffusive in $3D$ [11], diffusive in $2D$ and superdiffusive in $1D$ [12]. The long-range hopping $V(\mathbf{r})\propto r^{-2}$ is ubiquitous in pure two-dimensional systems [27], where it can be originated from the virtual exchange by two dimensional photons leading to the $2D$ dipole-dipole interaction [28] or indirect exchange by $2D$ electron-hole pairs leading to $2D$ RKKY interaction [29]. Power-law distant dependent hopping is crucial for the many-body localization problem [30; 31; 32; 33; 34; 35; 36; 37; 38; 39; 40; 41; 42; 43; 44; 45; 46; 47; 48; 49; 50; 51; 52; 53], where long-range interaction can result in localization breakdown at arbitrary disorder.
For the hopping under consideration $r^{-d}$ and weak disorder there is the transition in $3D$ to the standard delocalized phase characterized by the diffusive transport and ergodic dynamics [54], while in $1D$ eigenstates turns out to be multifractals with the dimension smaller than $1$ [12]. $2D$ systems are more complicated, because for the hopping $V(\mathbf{r})\propto r^{-d}$ the dimensionless Drude conductance diverges logarithmically with the system size as $C_{0}(L)=c_{*}\ln(L)$ [28]. Considering the balance of this logarithmic raise of conductance and its logarithmic suppression by coherent back scattering $c_{loc}\ln(L)$ Eq. (1), it was suggested in Ref. [28] that two delocalized phases can exist including the superdiffusive (fast) phase at $c_{*}>c_{loc}$ and the slow phase with diffusive transport at $c_{*}<c_{loc}$ and the phase boundary realized at $c_{*}=c_{loc}$. Yet it turns out that in systems, possessing time-reversal symmetry, for isotropic dipoles considered in Ref. [28] $c_{*}<c_{loc}$ for an arbitrarily disorder strength so the fast phase does not exist. This is in contrast with the systems with a broken time-reversal symmetry, possessing the transition between fast and slow phases, characterized by the unstable fixed point [28].
@@
Recent experimental realizations of $2D$ Anderson localization [56] and long-range hopping [57] represent the steps towards generating the settings targeted in the present work. Consequently, we believe that its experimental realization is possible and it is strongly encouraged.
@@
The phase boundary $c_{*}=c_{loc}$ is identified analytically in Secs. III, IV and verified numerically (see Sec. V and Fig. 1 there). Both phases are characterized by level statistics, that is of Wigner-Dyson for the fast phase and intermediate between Poisson and Wigner-Dyson otherwise (see Fig. 2), eigenstate dimension ($2$ for the fast phase and less than $2$ otherwise, as depicted in Fig. 3) and particle transport (superdiffusive or diffusive, see Fig. 4).
@@ II Model
$\widehat{H}=\sum_{i,j}V_{ij}c_{i}^{\dagger}c_{j}+\sum_{i}\phi_{i}c_{i}^{\dagger}c_{i},$ (2)
@@
For the Lorentzian distribution of random potentials the Green’s functions, averaged over the random potnetial realizations, can be evaluated exactly [58] in the momentum representation as
@@
where $V(\mathbf{q})$ expresses the Fourier transform of the hopping $V_{ik}$. For small wavevectors $q\ll 1$ it is given by the continuous approach that has the form
@@
as verified in Appendix A numerically, see also Ref. [20]. The Green functions are needed for the calculation of a zeroth order conductance given below. Below we set $V_{0}=1$ and use the effective disorder strength parameter $W$ to distinguish different phases.
III Zeroth order conductance
One can express the zeroth order dimensionless conductance tensor at given energy $E$ using the Kubo formula as [3; 7; 28]
$C^{(0)}_{ab}=\int\frac{d\mathbf{q}}{\pi(2\pi)^{d}}\left(\frac{\partial V(\mathbf{q})}{\partial q_{a}}\right)\left(\frac{\partial V(\mathbf{q})}{\partial q_{b}}\right)|{\rm Im}G(E,\mathbf{q})|^{2},$ (7)
@@ III Zeroth order conductance
The conductance in Eq. (7) diverges logarithmically at small wavevectors $q\rightarrow 0$ corresponding to the long distances. This divergence is caused by the divergence of the mean squared displacement $\int d^{2}\mathbf{r}V(r)^{2}r^{2}$ within the Fermi-Golden rule approach. At long distances the Fourier transform continuous representation becomes exact so the use of the hopping amplitude Fourier transform in the form of Eq. (6) is completely justified. Below we evaluate the diverging part needed for the characterization of the phase transition for the dipole - dipole interaction.
For the dipole-dipole hopping Eqs. (4), (6) the integral for the zeroth order conductance Eq. (7) diverges logarithmically at $q=0$. For the finite size $L$ the integral should be cutoff at the minimum value $q=\eta/L$, with the parameter $\eta=O(1)$ depending on the specific boundary conditions. Then the conductance tensor diverging components in Eq. (7) can be evaluated with the logarithmic accuracy as $C_{xx}=c_{x}\ln(L)$, $C_{yy}=c_{y}\ln(L)$, $c_{xy}=c_{yx}=0$, where
@@
The generalization to the arbitrary distribution of random potentials can be made by the replacements $E\rightarrow E-{\rm Re}\Sigma(E,0)$ and $W\rightarrow-{\rm Im}\Sigma(E,0)$ in Eqs. (8), (9), where $\Sigma(E,0)$ is the self-energy evaluated at energy $E$ and wavevector $q=0$. One should notice that $\Sigma(0,0)=0$. Finding the self-energy for arbitrary distribution of random potentials remains a challenge; yet this problem is much easier compared to the localization problem itself. The zeroth order conductance serves as an input parameter to the renormalization group equation for the conductance derived below.
@@
The initial conditions to Eq. (10) at large $p_{1}=O(1)$ are set using the zeroth order conductance as
@@
Similarly to the isotropic case Eq. (18) this equation has a stable fixed point at $c_{x}^{*}c_{y}^{*}<c_{loc}^{2}$. In the infinite size limit the steady state solution for conductance at that point can be approximated by
@@
Conductance approaches infinity in the infinite size limit for $c_{x}^{*}c_{y}^{*}\geq c_{loc}^{2}$. Consequently, the transition to the superdiffusive regime is defined as
@@
Figure 1: Phase diagram of the 2D interacting dipoles model and the color code shows the average level-spacing ratio $\langle r\rangle$, Eq. (26) characterizing the level statistics [60], evaluated for the system size $L=201$ and averaged over $1000$ realizations.
@@
For the dipole-dipole interaction the isotropic regime is realized only for the band center $E=0$, where the critical disorder is given by $W_{c}=\pi/\sqrt{(1+\sqrt{5})/2}\approx 2.47$. With decreasing disorder the fast phase emerges first at that energy. For the anisotropic regime realized at $E\neq 0$ the transition emerges at $c^{*}_{x}c^{*}_{y}=c_{loc}^{2}$, Eq. (25). Using the analytical results, Eq. (8), for the logarithmic growing rates $c^{*}_{x}$ and $c^{*}_{y}$ we determine the phase diagram depicted in Fig. 1, where solid lines indicate boundaries between slow and fast phases.
@@ V.1 Level Statistics
Energy level statistics is different for localized and delocalized states [62]. For delocalized states it approaches the Wigner-Dyson random matrix energy level statistics due to energy level repulsion, while energies of localized states are independent of each other and can be characterized by a Poisson statistics. Usually, Wigner-Dyson level statistics indicates ergodic behavior [60; 63], see, however, Ref.  [^1] Though there are cases when Wigner-Dyson [Kravtsov_2015] and Poisson [26; PhysRevLett.131.166401] level-spacing-ratio statistics coexists with the fractal extended eigenstates..
@@
In Fig. 2 we show the level statistics for the system of the size $L=291$ with the anisotropic dipole-dipole hopping Eq. (4) and different disorder strengths averaged over $200$ realizations.with the energy resolution $\delta E=0.1$. For the minimum disorder strength $W=1$, a substantial fraction of states with energies $|E|<2.7V_{0}$ indicated by the dashed line should belong to the fast phase (in all graphs we set $V_{0}=1$). The average level spacing ratio parameter $\langle r\rangle$ approaches the Wigner-Dyson limit $0.53$ in this domain as we expected for the fast phase. The intermediate disorder strength $W=2.5$ approximately corresponds to the last moment when the fast phase is present at $E=0$. For the strongest disorder $W=4$ all states suppose to belong to the slow phase. It is visually clear in Fig. 2 that our numerical findings are consistent with the assumption of ergodic behavior in the fast phase and its lack in the slow phase. The data for the level statistics are also presented in the phase diagram Fig. 1. They are consistent with the analytical results shown by the solid line.
@@ V.2 Fractal dimension
We define the fractal dimension using the informational dimension $D_{1}$ [65; 66] that can be expressed in terms of the average eigenstate wavefunction Shannon entropy $\xi(L)=-\langle\sum_{i}|c_{i}|^{2}\ln(|c_{i}|^{2})\rangle$, where the amplitudes $c_{i}$ represent eigenstate coefficients of the problem in the real-coordinate basis of sites $i$ with the energy $E$ close to zero ($-0.1<E<0.1$) and averaging is performed over all such states and different realizations of random potentials. Then the informational dimension is defined as $D_{1}=d\xi/d\ln(L)$, cf. Ref. [67]. We used this definition for numerical estimates of size-dependent fractal dimension calculating the functions $\xi$ for the sequence of lengths $L_{1}$, $L_{2}$, … $L_{n}$ arranged in ascending order and then numerically differentiating them. This yields $n-1$ estimates for fractal dimensions $D_{1}(l_{k})=(\xi(L_{k+1})-\xi(L_{k}))/\ln(L_{k+1}/L_{k})$ assigned to geometrically average sizes $l_{k}=\sqrt{L_{k}L_{k+1}}$ for $k=1,2...n-1$.
Numerical results should be compared with analytical estimates for fractal dimensions obtained using the generalized theory Eq. (17) and the connection between the dimension and the system conductance. established within the non-linear sigma model in Refs. [68; 69; 70; 8]. It was shown there that the informational dimension $D_{1}$ of a two-dimensional system with a finite conductance is smaller than the system dimension $2$. The difference of dimensions is inversely proportional to the dimensionless conductance at large conductances. Consequently, the fractal (informational) dimension can be expressed as $D_{1}=2-\eta_{d}c_{loc}/C$ at $C\gg c_{loc}$. Theory suggests [70; 8] $\eta_{d}=1$ for the non-linear sigma model with a short-range hopping.
@@
Here $C_{\infty}=c_{*}c_{loc}/(c_{loc}-c_{*})$ represents the infinite size limit of conductance in the slow phase $c_{*}<c_{loc}$ and $L_{0}$ is the unknown integration constant. We define this constant for each line shown in Fig. 3 minimizing the deviation of analytical ($2-1.3c_{loc}/C$) and numerical estimates of fractal dimensions. at largest sizes where the theory is most relevant.
@@
Figure 3: Finite size scaling of the fractal dimension vs. the predictions of the modified scaling theory of localization (see the text). The number of realizations is $40000$ for the minimum size $L=11$ decreasing with increasing the size to $2000$ for the maximum size $L=251$.

### v4

@@ Abstract
This universal scaling is limited to systems with a short-range hopping, while the hopping decreasing with the distance as $V(\mathbf{r})\propto r^{-d}$ or slower leads to delocalization of all states [1; 10; 11; 12; 13; 14] except for the marginal case of diverging Fourier transform of a hopping amplitude [15; 16; 17; 18; 19; 20; 21; 22; 23; 24; 25; 26]. If disorder is strong, eigenstates of the problem with the long-range hopping $V(\mathbf{r})\propto r^{-d}$ possess a multifractal structure and the time dependent displacement of the particle $r$ obeys the law $r\propto t^{1/d}$, which is subdiffusive in $3D$ [11], diffusive in $2D$ and superdiffusive in $1D$ [12]. The long-range hopping $V(\mathbf{r})\propto r^{-2}$ is ubiquitous in pure two-dimensional systems [27], where it can be originated from the virtual exchange by two-dimensional photons leading to the $2D$ dipole-dipole interaction [28] or indirect exchange by $2D$ electron-hole pairs leading to $2D$ RKKY interaction [29]. Power-law distant dependent hopping is crucial for the many-body localization problem [30; 31; 32; 33; 34; 35; 36; 37; 38; 39; 40; 41; 42; 43; 44; 45; 46; 47; 48; 49; 50; 51; 52; 53], where long-range interaction can result in localization breakdown at arbitrary disorder.
For the hopping under consideration $r^{-d}$ and weak disorder, there is the transition in $3D$ to the standard delocalized phase characterized by the diffusive transport and ergodic dynamics [54], while in $1D$ eigenstates turns out to be multifractals with the dimension smaller than $1$ [12]. $2D$ systems are more complicated, because for the hopping $V(\mathbf{r})\propto r^{-d}$ the dimensionless classical Drude conductance diverges logarithmically with the system size as $C_{0}(L)=c_{*}\ln(L)$ [28]. Considering the balance of this logarithmic raise of conductance and its logarithmic suppression by coherent back scattering $c_{loc}\ln(L)$ Eq. (1), it was suggested in Ref. [28] that two delocalized phases can exist including the superdiffusive (fast) phase at $c_{*}>c_{loc}$ and the slow phase with diffusive transport at $c_{*}<c_{loc}$ and the phase boundary realized at $c_{*}=c_{loc}$. Yet it turns out that in systems, possessing time-reversal symmetry, for isotropic dipoles considered in Ref. [28] $c_{*}<c_{loc}$ for an arbitrarily disorder strength so the fast phase does not exist. This is in contrast with the systems with a broken time-reversal symmetry, possessing the transition between fast and slow phases, characterized by the unstable fixed point [28].
@@
Recent experimental realizations of $2D$ Anderson localization [56] and long-range hopping [57] represent the steps towards generating the settings targeted in the present work. Consequently, we believe that its experimental realization is possible and it is strongly encouraged. Two dimensional $r^{-2}$ dipole-dipole interaction emerges in high dielectric constant films in a limited distance range where superdiffusive behavior can be observed as discussed in the end of Sec. IV.
In addition to the dipole-dipole interaction, there exists a long-range elastic $r^{-2}$ interaction in isolated dielectric films with a similar angular dependence to that for a dipole-dipole interaction. Similar phase transition is expected for that interaction without constraints like for a dipole-dipole interaction because elastic field is located fully inside the material. This is also true for an indirect exchange RKKY interaction within two dimensional metals. Of course for all interactions the consideration is limited to a subwavelength transport [31].
@@
The phase boundary $c_{*}=c_{loc}$ and a finite-size scaling of conductance are identified analytically in Secs. III, IV. The numerical study of Sec. V is targeted to check a consistency of the analytical theory with the results obtained by means of exact numerical diagonalization. Our investigation of level statistics in Sec. V.1 shows the consistency of the analytically predicted phase boundary with the behavior of the level statistics changing from the Wigner-Dyson statistics in the fast phase to the intermediate one between Poisson and Wigner-Dyson otherwise. The finite-size scaling of an eigenstate fractal dimensions reported in Sec. V.2 is consistent with the analytical theory under the assumption of a linear dependence of a fractal dimension on the inverse conductance, though the proportionality coefficient differs from the earlier theoretical predictions. Finally, we demonstrate that a time-dependent displacement of a particle initially localized in a single site shows super diffusive behavior in accord with the theoretical expectations, although the direct comparison of analytical and numerical results is problematic because of an insufficient maximum size of the system and the contribution of all eigenstates to the transport including the “slow” ones in the spectrum tails. All numerical results would not be fully conclusive for the infinite size limit without the analytical theory, which is the main outcome of the present work.
@@ II Model
$\widehat{H}=\frac{1}{2}\sum_{i,j}V_{ij}c_{i}^{\dagger}c_{j}+\sum_{i}\phi_{i}c_{i}^{\dagger}c_{i},$ (2)
@@
The present model is different from that of Ref. [28] because all transition dipole moments are oriented along the $x$ axis. To reproduce the settings of Ref. [28] we need to consider two degenerate states with identical random potentials in each site with perpendicular transition dipole moments and introduce the dipolar hopping between them accordingly. Also a short range hopping can be added between states with identical transition dipole moments. As noticed in the introduction in the present model there is the transition between fast and slow phases while it lacks in the isotropic model of Ref. [28].
For the Lorentzian distribution of random potentials the Green’s functions, averaged over the random potential realizations, can be evaluated exactly [58] in the momentum representation because averaging of local Green functions $1/(E-\phi_{i}-i\delta)$ with the distribution Eq. (3) yields $1/(E-iW)$ so averaging replaces all random potentials with the imaginary constant $iW$. Then the configurationally averaged Green function takes the form
@@
where $V(\mathbf{q})$ expresses the Fourier transform of the hopping $V_{ik}$. For small wavevectors $q\ll 1$ it can be approximated by the continuous limit of the dipole-dipole interaction Fourier transform with subtracted self-interaction as
@@
as verified in Appendix A numerically, see also Ref. [20]. The Green functions are needed for the calculation of a classical conductance given below. There we set $V_{0}=1$ and use the effective disorder strength parameter $W$ to distinguish different phases.
III Classical conductance
Long-range hopping results in a logarithmic divergence of a classical conductance. To characterize this divergence, let’s consider the generalized definition of a wavevector-dependent conductance [3; 6; 28] that is a target of the renormalization group analysis. It reads
[…]
where $g(\mathbf{p},\mathbf{p}_{1})$ stands for the Green functions taken at the energy of interest $E$ before the configurational averaging delimited-⟨$\langle$..$\rangle$ … ⟩.
Figure 1: Classical (A) and quantum (B) and (C) contributions to the dimensionless conductance.
The conductance can be expressed using diagrams [3; 6; 60] shown in Fig. 1, where solid lines stand for configurationally averaged Green functions Eq. (5) and dashed lines indicate correlations of two Green functions appearing due to the contributions of identical sites to both Green function. This series represents the conductance expansion in the ratio of the wavelength and the mean free path thus giving the quantum corrections Fig. 1.B to the classical conductance shown in Fig. 1.A. For the classical conductance input and output wavevectors are identical $\mathbf{p}=\mathbf{p}_{1}$. This classical conductance has a logarithmic divergence due to the squared group velocity terms like $\frac{\partial V(\mathbf{p})}{\partial p_{a}}^{2}\propto p^{-2}$, Eq. (6) lacking for the short-range hopping.
For the quantum corrections to the conductance shown in Fig. 1.B $\mathbf{p}\neq\mathbf{p}_{1}$ and, therefore, no logarithmic divergence emerges. Consequently, the correlations between two Green functions are neglected, when defining a positive logarithmically diverging contribution to the conductance, which we are interested in. This logarithmically diverging part is evaluated below exactly since in the present model exact Green functions are known Eq. (5).
The classical dimensionless conductance tensor at given energy $E$ takes the form (see Eq. (7) and Refs. [3; 7; 28])
$C^{(0)}_{ab}(\mathbf{q})=\int\frac{d\mathbf{p}}{\pi(2\pi)^{d}}\left(\frac{\partial V(\mathbf{p}+\mathbf{q}/2)}{\partial p_{a}}\right)\left(\frac{\partial V(\mathbf{p}-\mathbf{q}/2)}{\partial p_{b}}\right){\rm Im}G(E,\mathbf{p}+\mathbf{q}/2){\rm Im}G(E,\mathbf{p}-\mathbf{q}/2),$ (8)
@@ III Classical conductance
The conductance in Eq. (8) diverges logarithmically at small wavevectors $p$ corresponding to long distances. This divergence is caused by the divergence of the mean squared displacement $\int d^{2}\mathbf{r}V(r)^{2}r^{2}$ within the Fermi-Golden rule approach. At long distances the Fourier transform continuous representation becomes exact so the use of the hopping amplitude Fourier transform in the form of Eq. (6) is completely justified. Below we evaluate analytically the diverging part needed for the characterization of the phase transition for the anisotropic dipole - dipole hopping.
For the dipole-dipole hopping Eqs. (4), (6) the integral for the classical conductance Eq. (8) diverges logarithmically at $p=0$. Since this divergence takes place within the domain $q<p<1$ (for meaningful wavevectors $q>1/L$) the conductance tensor diverging components in Eq. (8) can be evaluated with the logarithmic accuracy as $C_{xx}=c_{x}\ln(1/q)$, $C_{yy}=c_{y}\ln(1/q)$, $c_{xy}=c_{yx}=0$, where
@@
The generalization to the arbitrary distribution of random potentials can be made by the replacements $E\rightarrow E-{\rm Re}\Sigma(E,0)$ and $W\rightarrow-{\rm Im}\Sigma(E,0)$ in Eqs. (9), (10), where $\Sigma(E,0)$ is the self-energy evaluated at energy $E$ and wavevector $q=0$. One should notice that $\Sigma(0,0)=0$. Finding the self-energy for arbitrary distribution of random potentials remains a challenge; yet this problem is much easier compared to the localization problem itself. The classical conductance serves as an input to the renormalization group equation for the conductance derived below in Sec. IV.
@@
The initial conditions to Eq. (11) at large $p_{1}=O(1)$ are set using the classical order conductance as
@@
Similarly to the isotropic case Eq. (19) this equation has a stable fixed point at $c_{x}^{*}c_{y}^{*}<c_{loc}^{2}$. In the infinite-size limit the steady state solution for conductance at that point can be approximated by
@@
Conductance approaches infinity in the infinite-size limit for $c_{x}^{*}c_{y}^{*}\geq c_{loc}^{2}$. Consequently, the transition to the superdiffusive regime is defined as
@@
Figure 2: Phase diagram of the 2D interacting dipoles model and the color code shows the average level-spacing ratio $\langle r\rangle$, Eq. (23) characterizing the level statistics [61], evaluated for the system size $L=201$ and averaged over $1000$ realizations separately for each value of $W$.
The solid line shows the analytical predictions for the phase boundary in the infinite system determined using Eq. (22).
@@
For the dipole-dipole interaction the isotropic regime is realized only for the band center $E=0$, where the critical disorder is given by $W_{c}=\pi/\sqrt{(1+\sqrt{5})/2}\approx 2.47$. With decreasing disorder the fast phase emerges first at that energy. For the anisotropic regime realized at $E\neq 0$ the transition emerges at $c^{*}_{x}c^{*}_{y}=c_{loc}^{2}$, Eq. (22). Using the analytical results, Eq. (9), for the logarithmic growing rates $c^{*}_{x}$ and $c^{*}_{y}$ we determine the phase diagram depicted in Fig. 2, where solid lines indicate analytical predictions for the boundary between slow and fast phases determined using Eq. (22).
The dipole-dipole hopping distance dependence $r^{-2}$ emerges in high-dielectric-constant films. For the film of the thickness $h$ possessing a dielectric constant $\epsilon$ substantially exceeding that of the environment ($\epsilon_{env}$), $1/r^{2}$ hopping amplitude distance dependence emerges for distances $r$ belonging to the domain $h<r<L_{max}=h\epsilon/\epsilon_{env}$ [62; 63] If the system is in the fast phase the conductance will grow within this domain reaching its maximum $C(L_{max})\approx(c_{*}-c_{loc})\ln(\epsilon/\epsilon_{env})$. At longer distances $L>L_{max}$ there is no long-range hopping contribution so the renormalization group equation takes the form $dC/d\ln(L)=-c_{loc}$ as for the short-range hopping. Then at distances exceeding $L_{max}$ a weak localization behavior of conductance is expected $C(L)=C(L_{max})-c_{loc}\ln(L/L_{max})$ until reaching the length $l\approx h(\epsilon/\epsilon_{env})^{\frac{c_{*}}{c_{loc}}}$ where $C(L)=0$. At that scale two dimensional Anderson localization is expected, so the size $l$ determines the localization length. It seems to be an exciting experimental challenge to investigate the excitation displacements vs. time in high-dielectric-constant films to observe all three regimes. For the hopping associated with elastic or RKKY interactions in isolated films there is no constraints like that for the dipole-dipole interaction so the superdiffusive behavior can be seen there at longer distances.
@@ V.1 Level Statistics
Energy level statistics is different for localized and delocalized states [65]. For delocalized states it approaches the Wigner-Dyson random matrix energy level statistics due to energy level repulsion, while energies of localized states are independent of each other and can be characterized by a Poisson statistics. Usually, Wigner-Dyson level statistics indicates ergodic behavior [61; 66], see, however, Ref.  [^1] Though there are cases when Wigner-Dyson [76] and Poisson [26; 77] level-spacing-ratio statistics coexists with the fractal extended eigenstates..
@@
In Fig. 3 we show the level statistics for the system of the size $L=291$ with the anisotropic dipole-dipole hopping Eq. (4) and different disorder strengths averaged over $200$ realizations.with the energy resolution $\delta E=0.1$. For the minimum disorder strength $W=1$, a substantial fraction of states with energies $|E|<2.7V_{0}$ indicated by the dashed line should belong to the fast phase (in all graphs we set $V_{0}=1$). The average level spacing ratio parameter $\langle r\rangle$ approaches the Wigner-Dyson limit $0.53$ in this domain as we expected for the fast phase. The intermediate disorder strength $W=2.5$ approximately corresponds to the last moment when the fast phase is present at $E=0$. For the strongest disorder $W=4$ all states suppose to belong to the slow phase. It is visually clear in Fig. 3 that our numerical findings are consistent with the assumption of ergodic behavior in the fast phase and its lack in the slow phase. The data for the level statistics are also presented in the phase diagram Fig. 2. They are consistent with the analytical results shown by the solid line. The results are shown for the maximum probed system size $L=251$. The results for smaller sizes are quite similar to those in Fig. 2 so we do not see any remarkable scaling of the level spacing ratio similarly to the earlier work [20]. This is in contrast with the fractal dimension scaling reported below.
@@ V.2 Fractal dimension
The numerical investigation of a fractal dimension reported in this section is targeted to verify a finite-size scaling of a conductance. predicted by the renormalization group theory Eq. (18) taking the advantage of the earlier predicted connection of conductance and dimension [68]. We define the fractal dimension using the informational dimension $D_{1}$ [69; 70] that can be expressed in terms of the average eigenstate wavefunction Shannon entropy $\xi(L)=-\langle\sum_{i}|c_{i}|^{2}\ln(|c_{i}|^{2})\rangle$, where the amplitudes $c_{i}$ represent eigenstate coefficients of the problem in the real-coordinate basis of sites $i$ with the energy $E$ close to zero ($-0.1<E<0.1$) and averaging is performed over all such states and different realizations of random potentials.
The informational dimension is connected to the fractal dimension $D_{q}$ of the multifractal eigenstate defined for the specific exponent $q$ as [20; 71]
$\langle\sum_{i}|c_{i}|^{2q}\rangle\propto L^{(1-q)D_{q}}.$ (24)
It corresponds to the limit of $q\rightarrow 1$, where the geometric averaging emerges naturally after both sides expansion in $1-q$ [71]. In this limit the fractal dimension is less sensitive to fluctuations thus reflecting a typical wavefunction behavior, which is our target.
For numerical calculations we define a size-dependent informational dimension as $D_{1}=d\xi/d\ln(L)$, cf. Ref. [72]. The dimension is estimated calculating the functions $\xi$ for the sequence of lengths $L_{1}$, $L_{2}$, … $L_{n}$ arranged in ascending order and then numerically differentiating them. This yields $n-1$ estimates for fractal dimensions $D_{1}(l_{k})=(\xi(L_{k+1})-\xi(L_{k}))/\ln(L_{k+1}/L_{k})$ assigned to geometrically average sizes $l_{k}=\sqrt{L_{k}L_{k+1}}$ for $k=1,2...n-1$.
Numerical results should be compared with analytical estimates for fractal dimensions obtained using the generalized theory Eq. (18) and the connection between the dimension and the system conductance. established within the non-linear sigma model in Refs. [73; 60; 68; 8]. It was shown there that the informational dimension $D_{1}$ of a two-dimensional system with a finite conductance is smaller than the system dimension $2$. The difference of dimensions is inversely proportional to the dimensionless conductance at large conductance $C\gg c_{loc}$. Consequently, the fractal (informational) dimension can be expressed as $D_{1}=2-\eta_{d}c_{loc}/C$ at $C\gg c_{loc}$. Theory suggests [68; 8] $\eta_{d}=1$ for the non-linear sigma model with a short-range hopping.
@@
Here $C_{\infty}=c_{*}c_{loc}/(c_{loc}-c_{*})$ represents the infinite-size limit of conductance in the slow phase $c_{*}<c_{loc}$ and $L_{0}$ is the unknown integration constant. We define this constant for each line shown in Fig. 4 minimizing the deviation of analytical ($2-1.3c_{loc}/C$) and numerical estimates of fractal dimensions. at largest sizes where the theory is most relevant.
@@
Figure 4: Finite-size scaling of the fractal dimension vs. the predictions of the modified scaling theory of localization (see the text). The number of realizations is $40000$ for the minimum size $L=11$ decreasing with increasing the size to $2000$ for the maximum size $L=251$.